OpenStudy (anonymous):

im so confused on how to do this 5/8c-8=12

OpenStudy (anonymous):

well you want to isolate C and that is the entire point on how to solve for c so your first step would be to get rid of one of those numbers and I would say to start with the -8 because its easier so do you know how to start?

OpenStudy (anonymous):

i have to add 8 to 12 i know that because adding 8 to -8 would make it 0

OpenStudy (anonymous):

yup so our new formula would look like this 5/8 C=20

OpenStudy (anonymous):

and that is the fraction 5/8 times C so you want to divide 20 by 5/8s so you would get C= 32 I believe :)

OpenStudy (anonymous):

how do you divide 5/8 into 20? thats the part i was getting stuck on

OpenStudy (anonymous):

the point is that the 5 is the 20 and the 8 is the 32 so for example lets say we have 2/4C=16 you say okay what times 4 gives me 16 and that would be 4 so then you have to do the same thing to both sides so 2 times 4 is 8 so the answer 8/16 so C for that formula is 8 +16 or 24 :)

OpenStudy (anonymous):

sorry its been a while and I am a bit tired lol hope that helps

OpenStudy (anonymous):

yes that helped lol thank youuuuuuu

OpenStudy (anonymous):

no problem